# Towards Automated Identification of Violation Symptoms of Architecture Erosion

### v4

@@ 3. Study Design
This section describes the research questions that motivated this research, followed by a detailed explanation of the data collection and the ML and DL models to automatically identify architecture violations from textual artifacts. Finally, we provide details of the validation survey and interview conducted to assess the effectiveness of automated identification of violation symptoms in practice.
Figure 2 provides an overview of the whole process. As illustrated in this figure (see the top box, Dataset Creation), we collected data and established a dataset on architecture violations in our previous work (Li2023wvs, ). During Phase 1, we explored the possibility of using this dataset to train ML/DL-based classifiers to automatically identify violation symptoms in code review comments. During Phase 2, we sent the survey questionnaire to 200 involved participants and received 24 responses that were used to validate the usefulness of automated identification of violation symptoms in practice.
@@ 3.1. Research Questions
In this work, we aim to explore the feasibility of automatically identifying violation symptoms of architecture erosion from code review comments. Specifically, we formulated the goal of this study by following the Goal-Question-Metric method (Basili1994gmq, ): analyze code review comments for the purpose of exploring the feasibility of automatically identifying violation symptoms of architecture erosion with respect to the performance on architecture erosion identification from the point of view of developers in the context of open source software development. To achieve our goal, we formulate three Research Questions (RQs):
RQ1: What is the performance of the classifiers in identifying violation symptoms from code review comments? RQ1.1: Which classifier performs best in identifying the violation symptoms? RQ1.2: Which word embedding models can help to improve the performance of classifiers? RQ1.3: To what extent, does the dimensionality of word embedding models impact the performance of classifiers?
Rationale: This RQ focuses on training feasible classifiers to automatically identify violation symptoms from textual artifacts. As mentioned in Section 1, the percentage of violation symptoms (belonging to architecture-level issues) is relatively lower than code-level issues. We thus expect that the description of violation symptoms is scattered in textual artifacts during development. RQ1.1 focuses on finding out the classifier with the best performance to distinguish violation symptoms from natural language by comparing several popular ML and DL algorithms. RQ1.2 aims at employing several well-known word embedding models to generate word vectors as features of labeled data, in order to choose the word embedding models that can help to improve the accuracy of classifiers. The word embedding models are a common ground for both ML and DL-based classifiers in our study. Finally, with RQ1.3, we seek to compare the differences in dimension size among different word embedding models, as it is one of the most important hyperparameters affecting the model performance(hutter2019aml, ). Specifically, RQ1.3 endeavors to shed light on the extent to which the selection of dimensionality impacts the performance of the classifiers in automatically identifying violation symptoms.
@@
Rationale: After training individual classifiers based on ML and DL algorithms, we want to know whether, and to what extent, an ensemble classifier composed of trained classifiers can help to improve the performance of identifying violation symptoms from code review comments. To answer this RQ, we compared the performance of individual classifiers with that of a combined classifier. More specifically, we employed a voting strategy to integrate the prediction values of each classifier and generated results after voting among individual classifiers. Answering this RQ can help to explore the possibility of further improving the performance of automated identification of violation symptoms.
@@
Rationale: If the results generated by the classifiers based on ML and DL algorithms are acceptable regarding the effectiveness of identified violation symptoms, automatically identifying violation symptoms from textual artifacts would be helpful in practice. Such automated techniques could facilitate tasks such as locating architecture violations and inspiring further investigation and remediation of similar architectural issues. To answer this RQ, we conducted a survey and semi-structured interviews to investigate how practitioners (both developers and reviewers) assess the effectiveness and applicability of the trained classifiers.
@@ 3.2. Data Collection
We relied on the dataset that we mined and manually identified in our previous study (Li2023wvs, ). As shown in Figure 2 (see Dataset Creation), we followed five steps to establish a specific dataset on architecture violations mined from code review comments. More specifically, we mined code review comments of four OSS projects in the OpenStack (i.e., Nova and Neutron) and Qt (i.e., Qt Base and Qt Creator) communities between 2014 and 2020. We mined the data through the REST API [^4] https://gerrit-review.googlesource.com/Documentation/rest-api.html supported by Gerrit. We first conducted experiment-based keyword refinement to improve our keyword search, as our previous experience indicated that a keyword-based approach could only capture very small proportions of code review comments containing violation symptoms (Li2023wvs, ). Subsequently, we proceeded with manual identification, and pilot and formal data labeling of review comments related to architecture violations. In total, we manually identified and labeled 606 code review comments pertaining to architecture violations. Code review comments were created for patches, and comments were made by code reviewers and developers. In terms of the data labeling process, four researchers first conducted a pilot data labeling by randomly selecting 50 review comments, to reach a consensus (the inter-rater agreement between the researchers measured in the Cohen’s Kappa coefficient value (Cohen1960cans, ) was 0.857) and ensure that we have the same understanding of violation symptoms. Any disagreements were discussed between the four researchers to reach a consensus. After that, the four researchers started the formal data labeling by dividing the retrieved 21,583 code review comments into four parts (each researcher manually labeled around 5,400 comments). After the formal data labeling, the first author checked the data labeling results from the other three researchers to make sure that there were no false positives. To mitigate potential bias, the four researchers discussed all the conflicts in the labeling results until we reached an agreement. To summarize, all data labeling results were checked by at least two researchers and conflicts were discussed among four researchers. Subsequently, all the collected code review comments regarding architecture violations (i.e., 606 code review comments) from the four projects were combined into an integrated dataset used in this work. The dataset has been made available in a replication package (onlinepackage_TSE, ).
3.3. Phase 1 - Automatic Classification of Violation Symptoms
@@
In this section, we describe the experimental setup that we followed to evaluate the performance of the trained classifiers (see Figure 3). The steps in this setup correspond to the steps of Phase 1 in Figure 2. First, we pre-processed the collected code review comments to generate structured word sets, and then utilized word embedding techniques to encode the words and generate vectors for presenting the words in review comments. Subsequently, the generated vectors acted as the input to train the classifiers based on ML/DL models, in order to learn to classify review comments as violations and non-violations. The experimental environment is a computer equipped with Intel(R) Core(TM) i7-10510U CPU and 16GB RAM, running Windows 11 (64-bit). The next subsections elaborate on these steps for both the ML and DL models. More details of the experiments (such as hyperparameters and experimental environment) can be found in the replication package (onlinepackage_TSE, ).
@@ 3.3.1. Machine Learning Models
[…]
@@
  • word2vec: word2vec was developed in 2013 (Mikolov2013word2vec, ) and it takes words as input and provides a word embedding matrix that contains high-dimensional unique vectors to represent each distinct word in given training corpora. word2vec can generate word embeddings through skip-gram and Continuous Bag-of-Words (CBOW) models. In this study, we adopted the pre-trained word2vec model proposed by Efstathiou et al. (Efstathiou2018weso, ). This model was trained based on the corpus comprising over 15 GB of textual data collected from Stack Overflow posts, which contains a plethora of textual information in the software engineering domain.
  • fastText: fastText is an open-source and lightweight library proposed by Meta AI Research lab in 2017 (Bojanowski2017fastText, ) and it is capable of learning text representations and training text classifiers. Based on the fastText model, Meta published pre-trained word vectors with 300 dimensions for 157 languages constructed on Common Crawl and Wikipedia [^5] https://fasttext.cc/docs/en/crawl-vectors.html. In this study, we employed the pre-trained fastText model, which can be transformed into 100 and 200-dimensional models, in order to control variables and avoid the impact of dimensionality on classifiers’ performance.
@@
Model Training. Given the binary classification task in this study, we selected five common ML algorithms that are widely used in classification tasks (Han2022dm, ), including Support Vector Machine (SVM), Logistic Regression (LR), Decision Tree (DT), Bernoulli Naive Bayes (NB), and k-Nearest Neighbor (kNN). These classification algorithms are also commonly applied in many areas of software engineering studies (Yang2022pmse, ).
For the ML classifiers, we performed a grid search to tune the hyperparameters of each model and picked the best-performing ML models with specific hyperparameters on the validation set. Grid search, an exhaustive search technique, explores every possible combination within a predefined search space to identify the optimal configuration. More specifically, this method is widely employed in the realms of machine learning and optimization and it involves a systematic evaluation that examines all of the combinations of a set of candidate settings to find the best hyperparameter combination (hutter2019aml, ).
@@ 3.3.2. Deep Learning Models
TextCNN is a convolutional neural network model that is commonly used for text classification tasks. This model takes a sentence as input and applies multiple convolutional filters to extract important features from the text. A TextCNN classifier based on the vocabulary set of a dataset uses a one-hot encoding scheme to represent words and generate the word embedding, while a TextCNN classifier can also utilize pre-trained word embedding models to capture semantic information of words. Additionally, the TextCNN model requires a proper setting to construct neural network models, such as hyperparameters and word embedding parameters. In this work, we set the embedding dimension to 200, the learning rate to 0.001, and the dropout to 0.25 to handle the overfitting, while the filter region size is (3, 4, 5). Each DL model’s batch size is set to 16, and the network is trained for 100 epochs with the early stopping strategy with the purpose of avoiding overfitting. We configured the patience parameter to 8, indicating that the training process would terminate after 8 consecutive epochs with no observed improvement in model performance. These approaches are common practices in machine learning, aimed at preventing overfitting and optimizing the model’s convergence (Watson2022SLRDL, ).
@@ 3.3.3. Performance Evaluation
To avoid the class imbalance problem (he2009imbalance, ), a data distribution problem that significantly impacts classifiers’ performance, we constructed our dataset with a balanced class distribution (i.e., the same number of review comments labeled as “violations” and “non-violations”), which avoids the poor generalization ability of classifiers. With respect to the model training and validation strategy, we employed 10-fold cross-validation when we conducted a performance evaluation of the classifiers. We divided the collected dataset into three parts, i.e., we randomly set aside 80% of the data for the training set (60%) and validation set (20%), and 20% for the test set. Besides, we set a fixed seed for reproducibility during experiments.
3.4. Ensemble Classifier
To identify architecture violations from textual artifacts (i.e., code review comments), we trained several ML and DL-based classifiers to predict the labels of new review comments. However, the trained classifiers have varying performances in predicting the labels of target artifacts. In RQ2, we asked whether composing the individual classifiers into an ensemble classifier can help improve the overall performance. This composition can be done by utilizing the ensemble learning technique, i,e., majority voting strategy (Dietterich2000emml, ). Specifically, a combined classifier of several individual classifiers is generated with the purpose of building an improved classifier, which predicts an output (class) based on the highest probability of chosen class as the output (Dietterich2000emml, ). This aggregating criterion is like an election process and can produce combined voting results from each classifier’s output.
The voting criteria include Hard Voting (voting is calculated on the predicted output class) and Soft Voting (voting is calculated on the predicted probability of the output class). To reduce the training time, we employed the Hard Voting (a simple majority voting strategy) strategy in this work. In other words, the final predicted results depend on the majority output (i.e., labels) of the individual classifiers.
@@ 3.5. Phase 2 - Validation: Survey and Interview
(1) Survey. Surveys can be conducted in different ways, such as online questionnaires, phone surveys (Lethbridge2005sse, ). Our survey is an online questionnaire of the cross-sectional type, which allows us to collect data and compare differences between participants at a single point in time. We adopted self-administered questionnaires to collect responses, since (a) respondents can answer the survey at their convenience (reducing our time investment and increasing the response rate); (b) questionnaires can reduce research bias since the researcher does not have any contact with the respondents when they fill out the survey; and (c) respondents enjoy better privacy than interviews.
@@
We invited participants through customized emails (see the replication package (onlinepackage_TSE, )) attaching the links to their code review comments that contain violation symptoms and are identified by our approach (i.e., the best-performing classifier), as well as the survey form (also see the replication package (onlinepackage_TSE, )). Our survey was constructed based on four parts: Problem Statement, Our approach, Survey Goal, and Survey Questions. Problem Statement introduces the research background regarding what is architecture erosion and its violation symptoms, as well as why it is important to investigate violation symptoms. Our approach describes the approach that we developed to automatically identify violation symptoms with high performance from a large number of code reviews as well as two examples of violation symptoms of architecture erosion detected by our approach from code review comments. In Survey Goal, we explain the purpose of this survey. In Survey Questions, we ask five general questions and present six statements concerning the usefulness of our approach. Note that, before the formal survey, we did a pilot survey (receiving 6 responses) with the purpose of revising and improving our survey questions and examples.
@@
The five general survey questions include demographic information (e.g., their countries and work experience). The six statements are assessed using a five-point Likert scale (Strong Agree, Agree, Neutral, Disagree, and Strongly Disagree) with one option (“I don’t know”). This scale enables participants to rate the extent that they agree or disagree with the statements. Finally, at the end of the survey, we ask an open-ended question to let the respondents freely provide comments and suggestions about our survey. To formulate the statements, we took inspiration from the survey questions of two survey studies (Nasab2021ais, ; Khalajzadeh2022sda, ) that evaluate the usefulness of automatic classifiers based on ML and DL techniques in the software engineering community. We chose to use declarative statements in our validation survey since declarative statements are deemed appropriate for exploratory studies to convey simple and clear judgements; in contrast, descriptive statements are more suitable in explanatory studies to provide the details of the judgements in various aspects. The statements are as follows:
@@
(2) Interview. As a supplement to the survey, we also invited the same practitioners to participate in semi-structured interviews, which allow for more in-depth discussions. Specifically, the interviews focused on the developers’ opinions on the effectiveness of our trained classifiers, as well as the role of such automated techniques in the code review process during software development. The semi-structured interviews were conducted by focusing on the following three open-ended Interview Questions (IQs):
@@
  • IQ1: How do you think the integration of an automated system for identifying violation symptoms of architecture erosion from code review comments would fit within the code review workflow? This system would analyze reviewer comments in real-time to flag potential or overlooked violation symptoms. By doing so, it would prevent the incorporation of code changes that violate system architecture, thereby contributing to a more robust and reliable development process. What are your thoughts on the potential benefits and challenges of such integration?
@@
  • IQ2: Based on your experience, what strategies would you recommend to effectively use automated identification of violation symptoms in the code review process? How can this technology be utilized to maximize its benefits and efficiency in identifying and addressing architectural issues during code review?
@@
  • IQ3: In addition to code review scenarios, can you envision other contexts or scenarios where our approach for automated identification of violation symptoms could be beneficial in software development practice? What potential applications or integrations do you see as valuable in enhancing code quality and development workflows?
@@
Participants. Our collected data from Gerrit contains the developers’ information (i.e., names and email addresses); this information is not included in the replication package for privacy reasons. For the survey, we invited participants to fill out our survey by sending emails to the developers whose comments are identified by our classifiers. During the pilot survey, we got responses from 6 developers. The pilot survey helped us to evaluate and refine our survey design. For example, we added more precise examples of violation symptoms from code review comments in our survey in order to enhance practitioners’ understanding, and we refined the content of the customized emails based on the received feedback from the pilot survey participants. Then, we invited the remaining developers involved in the discussions on violation symptoms to fill in the survey. Finally, for the semi-structured interviews, we invited the above-mentioned participants to give feedback on our trained classifiers and the role of such automated techniques during code review. We first conducted two pilot interviews and found that, the interview questions are suitable. Subsequently, we interviewed the remaining participants who accepted our invitation, resulting in a total of six interviews.
@@
Sample and Population. The target population is limited to developers whose code review comments are identified by our approach. We used a Non-Probabilistic Sampling method in this survey, namely, Convenience Sampling. As a type of non-probability sampling method, Convenience Sampling can help us to conveniently reach a suitable number of developers who are willing to participate in our survey. The reason we chose this sampling method is that we could not intentionally choose a sample due to the limitations of geographical constraints and scheduling conflicts.

### v5

@@ 3. Study Design
This section describes the research questions that motivated this research, followed by a detailed explanation of the data collection and the ML/DL models to automatically identify architecture violations from textual artifacts. Then, we provide details of the validation survey and interview conducted to assess the effectiveness of automated identification of violation symptoms in practice. Furthermore, we introduce and compare the state-of-the-art LLMs on the same task with traditional ML/DL approaches.
Figure 2 provides an overview of the study design. As illustrated in this figure (see the top box, Dataset Creation), we collected data and established a dataset on architecture violations in our previous work (Li2023wvs, ), which contains code review comments related to architecture violations During Phase 1, we explored the possibility of using this dataset to train ML/DL-based classifiers to automatically identify violation symptoms in code review comments. During Phase 2, we sent the survey questionnaire to 200 involved participants and received 24 responses that were used to validate the usefulness of automated identification of violation symptoms in practice. During Phase 3, we constructed LLM-based classifiers through prompts using state-of-the-art LLMs and compared their performance against the ML/DL methods from Phase 1.
@@ 3.1. Research Questions
In this work, we aim to explore the feasibility of automatically identifying violation symptoms of architecture erosion from code review comments. Specifically, we formulated the goal of this study by following the Goal-Question-Metric method (Basili1994gmq, ): analyze code review comments for the purpose of exploring the feasibility of automatically identifying violation symptoms of architecture erosion with respect to the performance on architecture erosion identification from the point of view of developers in the context of open source software development. To achieve our goal, we formulate four Research Questions (RQs):
RQ1: How accurately can classifiers identify violation symptoms from code review comments? RQ1.1: Which classifier performs best in identifying the violation symptoms? RQ1.2: Which word embedding models can help to improve the performance of classifiers? RQ1.3: To what extent, does the dimensionality of word embedding models impact the performance of classifiers?
Rationale: This RQ focuses on training feasible classifiers to automatically identify violation symptoms from textual artifacts. The percentage of violation symptoms (belonging to architecture-level issues) is relatively lower than code-level issues. We thus expect that the description of violation symptoms is scattered in textual artifacts during development. RQ1.1 focuses on finding out the classifier with the best performance to distinguish violation symptoms from natural language by comparing several popular ML/DL algorithms. RQ1.2 aims at employing several well-known word embedding models to generate word vectors as features of labeled data, in order to choose the word embedding models that can help to improve the accuracy of classifiers. The word embedding models are a common ground for both ML/DL-based classifiers in our study. Finally, with RQ1.3, we seek to compare the differences in dimension size among different word embedding models, as it is one of the most important hyperparameters affecting the model performance(hutter2019aml, ). Specifically, RQ1.3 endeavors to shed light on the extent to which the selection of dimensionality impacts the performance of the classifiers in automatically identifying violation symptoms.
@@
Rationale: After training individual classifiers based on ML/DL algorithms, we want to know whether, and to what extent, an ensemble classifier composed of trained classifiers can help to improve the performance of identifying violation symptoms from code review comments. To answer this RQ, we compared the performance of individual classifiers with that of a combined classifier. More specifically, we employed a voting strategy to integrate the prediction values of each classifier and generated results after voting among individual classifiers. Answering this RQ can help to explore the possibility of further improving the performance of automated identification of violation symptoms.
@@
Rationale: If the results generated by the classifiers based on ML/DL algorithms are acceptable regarding the effectiveness of identified violation symptoms, automatically identifying violation symptoms from textual artifacts would be helpful in practice. Such automated techniques could facilitate tasks such as locating architecture violations and inspiring further investigation and remediation of similar architectural issues. To answer this RQ, we conducted a survey and semi-structured interviews to investigate how practitioners (both developers and reviewers) assess the effectiveness and applicability of the trained classifiers.
RQ4: Do state‑of‑the‑art LLM-based classifiers truly outperform traditional ML/DL classifiers on the same classification tasks, and if so, to what extent do they offer performance gains?
Rationale: Considering the remarkable performance of LLMs on various SE tasks (hou2024large, ; li2025ChatGPT, ), we aim to further validate our conjecture whether LLM-based classifiers truly outperform traditional ML/DL-based classifiers regarding the automated identification of violation symptoms in code review comments. To answer RQ4, we developed several LLM-based classifiers using different prompt strategies to perform the same classification tasks in RQ1 and RQ2, enabling a fair comparison of performance gains.
@@ 3.2. Data Collection
We relied on the dataset that we mined and manually identified in our previous study (Li2023wvs, ). As shown in Figure 2 (see Dataset Creation), we followed five steps to establish a specific dataset on architecture violations mined from code review comments. More specifically, we mined code review comments of four OSS projects in the OpenStack (i.e., Nova and Neutron) and Qt (i.e., Qt Base and Qt Creator) communities between 2014 and 2020. We mined the data through the REST API [^4] https://gerrit-review.googlesource.com/Documentation/rest-api.html supported by Gerrit. We first conducted experiment-based keyword refinement to improve our keyword search, as our previous experience indicated that a keyword-based approach could only capture very small proportions of code review comments containing violation symptoms (Li2023wvs, ). The dataset collection process comprised three main steps: keywords searching, semantically-similar words querying (identifying potentially semantic-related data), and co-occurrence matrix-based querying (capturing potentially co-occurred and associated data). The whole manual labeling and analysis process took the researchers around one and a half months. Subsequently, we proceeded with manual identification, and pilot and formal data labeling of review comments related to architecture violations. In total, we manually identified and labeled 606 code review comments pertaining to architecture violations. Code review comments were created for patches, and comments were made by code reviewers and developers. In terms of the data labeling process, four researchers first conducted a pilot data labeling by randomly selecting 50 review comments, to reach a consensus (the inter-rater agreement between the researchers measured in the Cohen’s Kappa coefficient value (Cohen1960cans, ) was 0.857) and ensure that we have the same understanding of violation symptoms. Any disagreements were discussed between the four researchers to reach a consensus. After that, the four researchers started the formal data labeling by dividing the retrieved 21,583 code review comments into four parts (each researcher manually labeled around 5,400 comments). After the formal data labeling, the first author checked the data labeling results from the other three researchers. Meanwhile, to mitigate potential bias, the four researchers discussed all the conflicts in the labeling results until we reached an agreement. To summarize, all data labeling results were checked by at least two researchers and conflicts were discussed among four researchers. Subsequently, all the collected code review comments regarding architecture violations (i.e., 606 code review comments) from the four projects were combined into an integrated dataset used in this work. The dataset is available in our replication package (onlinepackage, ).
3.3. Phase 1 - Automated Classification of Violation Symptoms based on ML/DL Techniques
@@
In this section, we describe the experimental setup that we followed to evaluate the performance of the trained classifiers (see Figure 3). The steps in this setup correspond to the steps of Phase 1 in Figure 2. First, we pre-processed the collected code review comments to generate structured word sets, and then utilized word embedding techniques to encode the words and generate vectors for presenting the words in review comments. Subsequently, the generated vectors acted as the input to train the classifiers based on ML/DL models, in order to learn to classify review comments as violations and non-violations. The experimental environment is a computer equipped with Intel(R) Core(TM) i7-10510U CPU and 16GB RAM, running Windows 11 (64-bit). Strictly speaking, DL is a subset of ML. However, to clearly differentiate between traditional ML and DL approaches, we provide a detailed explanation of these steps for ML/DL models separately in the following subsections. More details of the experiments (such as hyperparameters and experimental environment) can be found in the replication package (onlinepackage, ).
@@ 3.3.1. Machine Learning Models
Step 2. Noise Removal: Developers usually use plain text when they discuss violation symptoms during code review, which contain various noise data such as punctuation, numbers, and special characters (e.g., $\backslash$”, “*”). Noise data usually does not contain valuable semantic information, and we therefore removed it.
@@
  • word2vec: word2vec was developed in 2013 (Mikolov2013word2vec, ) and it takes words as input and provides a word embedding matrix that contains high-dimensional unique vectors to represent each distinct word in given training corpora. word2vec can generate word embeddings through skip-gram and Continuous Bag-of-Words (CBOW) models. In this study, we adopted the pre-trained word2vec model proposed by Efstathiou et al. (Efstathiou2018weso, ). This model was trained based on the corpus comprising over 15 GB of textual data collected from Stack Overflow posts, which contains a plethora of textual information in the SE domain.
  • fastText: fastText is an open-source and lightweight library proposed by Meta AI Research lab in 2017 (Bojanowski2017fastText, ) and it is capable of learning text representations and training text classifiers. Based on the fastText model, Meta published pre-trained word vectors with 300 dimensions for 157 languages constructed on Common Crawl and Wikipedia [^5] https://fasttext.cc/docs/en/crawl-vectors.html. In this study, we employed the pre-trained fastText model, which can be transformed into 100 and 200-dimensional models, to control variables and avoid the impact of dimensionality on classifiers’ performance.
@@
Model Training. Given the binary classification task in this study, we selected five common ML algorithms that are widely used in classification tasks (Han2022dm, ), including Support Vector Machine (SVM), Logistic Regression (LR), Decision Tree (DT), Bernoulli Naive Bayes (NB), and k-Nearest Neighbor (kNN). These algorithms are also commonly applied in areas of SE studies (Yang2022pmse, ).
For the ML classifiers, we performed a grid search to tune the hyperparameters of each model and picked the best-performing ML models with specific hyperparameters on the validation set. Grid search, an exhaustive search technique, explores every possible combination within a predefined search space to identify the optimal configuration. Specifically, it is widely employed in the realms of machine learning and optimization and it involves a systematic evaluation that examines all of the combinations of a set of candidate settings to find the best hyperparameter combination (hutter2019aml, ).
@@ 3.3.2. Deep Learning Models
TextCNN is a convolutional neural network model that is commonly used for text classification tasks. This model takes a sentence as input and applies multiple convolutional filters to extract important features from the text. A TextCNN classifier based on the vocabulary set of a dataset uses a one-hot encoding scheme to represent words and generate the word embedding, while a TextCNN classifier can also utilize pre-trained word embedding models to capture semantic information of words. Additionally, the TextCNN model requires a proper setting to construct neural network models, such as hyperparameters and word embedding parameters. In this work, we set the embedding dimension to 200, the learning rate to 0.001, and the dropout to 0.25 to handle the overfitting, while the filter region size is (3, 4, 5). Each DL model’s batch size is set to 16, and the network is trained for 100 epochs with the early stopping strategy with the purpose of avoiding overfitting. We configured the patience parameter to 8, indicating that the training process would terminate after 8 consecutive epochs with no observed improvement in model performance. These approaches and settings are common practices in work related to ML/DL and text classification, aimed at preventing overfitting and optimizing the model’s convergence (Watson2022SLRDL, ).
@@ 3.3.3. Performance Evaluation
To avoid the class imbalance problem (he2009imbalance, ; Buda2018SMSCNN, ), a data distribution problem that significantly impacts classifiers’ performance, we constructed a balanced dataset containing equal numbers of review comments labeled as “violation” and “non-violation”, respectively. Note that, we randomly selected 606 “non-violation” comments to balance the dataset, and this balanced distribution mitigates the poor generalization ability of classifiers, and avoids leading to highly misleading assessment (Buda2018SMSCNN, ). With respect to the model training and validation strategy, we employed 10-fold cross-validation when we conducted a performance evaluation of the classifiers. We divided the collected dataset into three parts, i.e., we randomly set aside 80% of the data for the training set (60%) and validation set (20%), and 20% for the test set. Besides, we set a fixed seed for reproducibility during experiments.
3.3.4. Ensemble Classifier
To identify architecture violations from textual artifacts (i.e., code review comments), we trained several ML/DL-based classifiers to predict the labels of new review comments. However, the trained classifiers have varying performances in predicting the labels of target artifacts. In RQ2, we asked whether composing the individual classifiers into an ensemble classifier can help improve the overall performance. This composition can be done by utilizing the ensemble learning technique, i,e., majority voting strategy (Dietterich2000emml, ). Specifically, a combined classifier of several individual classifiers is generated with the purpose of building an improved classifier, which predicts an output (class) based on the highest probability of chosen class as the output (Dietterich2000emml, ). This aggregating criterion is like an election process and can produce combined voting results from each classifier’s output.
The voting criteria include Hard Voting (voting is calculated on the predicted output class) and Soft Voting (voting is calculated on the predicted probability of the output class). Considering the computational and memory requirements, we employed the Hard Voting strategy (a cost-efficient majority voting strategy) in this work. In other words, the final predicted results depend on the majority output (i.e., labels) of the individual classifiers.
@@ 3.4. Phase 2 - Validation: Survey and Interview
(1) Survey. Surveys can be conducted in different ways, such as online questionnaires, phone surveys (Lethbridge2005sse, ). Our survey is an online questionnaire of the cross-sectional type, which allows us to collect data and compare differences between participants at a single point in time. We adopted self-administered questionnaires to collect responses, since (a) respondents can answer the survey at their convenience (reducing their time investment and increasing the response rate); (b) questionnaires can reduce research bias since the researcher does not have any contact with the respondents when they fill out the survey; and (c) respondents enjoy better privacy than interviews.
@@
We invited all the developers involved in our collected dataset through customized emails (see the replication package (onlinepackage, )) attaching the links to their code review comments that contain violation symptoms and are identified by our approach (i.e., the best-performing classifier), as well as the survey form (also see the replication package (onlinepackage, )). Our survey was constructed based on four parts: Problem Statement, Our approach, Survey Goal, and Survey Questions. Problem Statement introduces the research background regarding what is architecture erosion and its violation symptoms, as well as why it is important to investigate violation symptoms. Our approach describes the approach that we developed to automatically identify violation symptoms with high performance from a large number of code reviews as well as two examples of violation symptoms of architecture erosion detected by our approach from code review comments. In Survey Goal, we explain the purpose of this survey. In Survey Questions, we ask five general questions and present six statements concerning the usefulness of our approach. Note that, before the formal survey, we did a pilot survey (receiving 6 responses) with the purpose of revising and improving our survey questions and examples.
@@
The five general survey questions include demographic information (e.g., their countries and work experience). The six statements are assessed using a five-point Likert scale (Strong Agree, Agree, Neutral, Disagree, and Strongly Disagree) with one option (“I Don’t Know”). This scale enables participants to rate the extent that they agree or disagree with the statements. At the end of the survey, we ask an open-ended question to let the respondents freely provide comments and suggestions about our survey. To formulate the statements, we took inspiration from the survey questions of two survey studies (Nasab2021ais, ; Khalajzadeh2022sda, ) that evaluate the usefulness of automatic classifiers based on ML/DL techniques in the SE community. We chose to use declarative statements in our validation survey since declarative statements are deemed suitable for exploratory studies to convey simple and clear judgments; in contrast, descriptive statements are more suitable in explanatory studies to provide detailed judgments in various aspects. The statements are as follows:
@@
(2) Interview. As a supplement to the survey, we also invited the same practitioners to participate in semi-structured interviews for more in-depth discussions. Specifically, the interview focused on the developers’ opinions on classifier effectiveness, as well as the role of such automated techniques in the code review process during software development. The semi-structured interviews were conducted by the following three open-ended Interview Questions (IQs):
@@
  • IQ1: How do you think the integration of an automated system for identifying violation symptoms of architecture erosion from code review comments would fit within the code review workflow? This system would analyze reviewer comments in real-time to flag potential or overlooked violation symptoms. By doing so, it would prevent the incorporation of code changes that violate system architecture, thereby contributing to a more robust and reliable development process. What are your thoughts on the potential benefits and challenges of such integration? This question aims to assess the practical feasibility and acceptance of integrating automated violation symptoms identification into existing code review workflows.
@@
  • IQ2: Based on your experience, what strategies would you recommend to effectively use automated identification of violation symptoms in the code review process? How can this technology be utilized to maximize its benefits and efficiency in identifying and addressing architectural issues during code review? This question seeks to gather expert insights on optimal deployment strategies and best practices for automated violation symptoms identification.
@@
  • IQ3: In addition to code review scenarios, can you envision other contexts or scenarios where our approach for automated identification of violation symptoms could be beneficial in software development practice? What potential applications or integrations do you see as valuable in enhancing code quality and development workflows? This question explores the broader applicability and extensibility of automated violation symptoms identification beyond code review contexts.
@@
Participants. Our collected data from Gerrit contains the developers’ information (i.e., names and email addresses); this information is not included in the replication package for privacy reasons. For the survey, we invited the potential participants to fill out our survey by sending emails to all developers whose review comments are included in our collected dataset. During the pilot survey, we got responses from 6 developers. The pilot survey helped us to evaluate and refine our survey design. For example, we added more precise examples of violation symptoms from code review comments in our survey in order to enhance practitioners’ understanding, and we refined the content of the customized emails based on the received pilot feedback. Then, we invited the remaining developers involved in the discussions on violation symptoms to fill in the survey. Finally, for the semi-structured interviews, we invited the above-mentioned participants to give feedback on our trained classifiers and the role of such automated techniques during code review. We first conducted two pilot interviews and found that, the interview questions are suitable. Subsequently, we interviewed the remaining participants who accepted our invitation, resulting in a total of six interviews.
@@
Sample and Population. The target population is limited to developers whose code review comments are identified by our approach. We used a Non-Probabilistic Sampling method in this survey, namely, Convenience Sampling. As a type of non-probability sampling method, Convenience Sampling can help conveniently reach a suitable number of developers who are willing to participate in our survey. The reason we chose this sampling method is that we could not intentionally choose a sample due to the limitations of geographical constraints and scheduling conflicts.